OpenStudy (anonymous):

Challenging integral: Integrate from 0--> infinity the argument: (e^(-z*x^2)*x^(n+1)*J_n(c*x) dx)

OpenStudy (bahrom7893):

@Zarkon

OpenStudy (bahrom7893):

@TuringTest

OpenStudy (bahrom7893):

@joemath314159

OpenStudy (anonymous):

\[\Large \int_0^\infty e^{-zx^2}x^{n+1}J_ncx dx\]What is \(\large J_n\)? A constant?

OpenStudy (anonymous):

Oh wait, it's a function...? \[\Large \int_0^\infty e^{-zx^2}x^{n+1}J_n(cx) dx\]

OpenStudy (anonymous):

yea its the nth order bessel function

OpenStudy (anonymous):

I use the recursion relation: J_n = SUM ( (-1)^k / (gamma(k+1) *gamma(k+n+1)) *(1/2x)^(2k+n) where gamma is the gamma function. Just the definition of the J_n function. Not sure how to integrate it...

OpenStudy (cruffo):

Is it true that the Bessel functions are normalized so that \[\int_0^\infty J_n(x) dx=1 ?\]

OpenStudy (anonymous):

yup they are

OpenStudy (cruffo):

and you are using \[\large J_n(x) = \sum_{k=0}^\infty \frac{(-1)^k}{\Gamma(k+1)\Gamma(k+n+1)}\left( \frac{1}{2}x\right) ^{2k+n}\]

OpenStudy (anonymous):

yep

OpenStudy (cruffo):

how did you adjust for J_n(cx) ? Is is just \[\large J_n(cx) = \sum_{k=0}^\infty \frac{(-1)^k}{\Gamma(k+1)\Gamma(k+n+1)}\left( \frac{c}{2}x\right) ^{2k+n}\]

OpenStudy (anonymous):

I believe so yea

OpenStudy (cruffo):

Is this from QM or probability ???

OpenStudy (anonymous):

electromagnatism

OpenStudy (cruffo):

I don't claim to know anything about EM. Usually there is some trick when working with distributions of this type, so that you don't have to go through all the integration mess. However, have you tried this approach (see pdf attached)

OpenStudy (cruffo):

About the messy repeated integration by parts see http://www.colby.edu/chemistry/PChem/notes/Integral.pdf

OpenStudy (cruffo):

humm....

OpenStudy (anonymous):

Thanks that helped. I did a bit of re-shuffling of some terms and got the right answer.

OpenStudy (cruffo):

cool!